Project

General

Profile

Actions

Physics Shift Report

Start Date: 22 June
End Date: 26 June
Start Run: 7076
End Run: 7152
Physics shifter: Paolo Franchini

Summary

Several times the CDB values are lower than the DAQ/run status, i.e.:

run 7078
              cdb    daq
ToF0 Triggers 77149    77168
ToF1 Triggers 22176    21182
ToF2 Triggers 4448    4450

run 7081
              cdb    daq
ToF0 Triggers 76926    77001
ToF1 Triggers 20818    20842
ToF2 Triggers 4809    4814

run 7082
              cdb    daq
ToF0 Triggers 79831    79918
ToF1 Triggers 21524    21544
ToF2 Triggers 4175    4182

run 7083
              cdb    daq
ToF0 Triggers 80510    80599    
ToF1 Triggers 21855    21883
ToF2 Triggers 3919    3920

Run documentation

  • Pion reference beam
  • 1000 spills of each diffuser setting
  • 190 MeV pion beam
  • 290 MeV pion beam
  • 6,200 M0 beam
  • 240 MeV muon beam

Elog

Activity well documented.

Run Condition Summary Spreadsheet/CDB

Always updated and still necessary (Pierrick-Janusz standing action). The last one is in http://micewww.pp.rl.ac.uk/attachments/download/4119

Analysis

No problems in the data movement. Minor issues solved by the day after.
All the plots produced by the physics devil are empty.

Updated by Franchini, Paolo almost 8 years ago ยท 11 revisions